OpenStudy (anonymous):

limit as x approaches -infinity x / (x - (sqrt. x^2 + 7))

OpenStudy (anonymous):

How many ways can u stack 77 books into equal pile?

OpenStudy (anonymous):

i don't know? how does this help

OpenStudy (anonymous):

\[\lim_{x\rightarrow \infty} \frac{x}{x- \sqrt{x^2+7}}= \lim_{x\rightarrow \infty} \frac{x}{x}\cdot \frac{1}{1-\sqrt{1+7/x^2}} \] does that help?

OpenStudy (anonymous):

that does. but then you'd get 1/0 that DNE

OpenStudy (anonymous):

DNE is not the answer

OpenStudy (anonymous):

I'm afraid the limit does not exist.

OpenStudy (anonymous):

Unless you refer to the limit as negative infinity.